# Scaling Limits of Line Models in Degenerate Environment

### v1

@@ Abstract
We consider a 2-dimensional model of random walk in random environment known as line model. The environment is described by two independent families of i.i.d. random variables dictating rates of jumps in vertical, respectively horizontal directions, and whose values are constant along vertical, respect. horizontal lines. When jump rates are heavy-tailed in one of the directions, we prove that the random walk becomes superdiffusive in that direction, with an explicit scaling limit written as a two-dimensional Brownian motion time-changed (in one of the components) by a Kesten-Spitzer process as introduced in [19]. In the case of a fully degenerate environment, a sufficient condition for non-explosion is provided, and conjectures on the associated scaling limit are formulated.
@@ 1.1 Outline of the article
This paper deals with a time-continuous random walk $(X(t))_{t\geq 0}$ on $\mathbb{Z}^{2}$ in the random environment $\omega$. The walk to nearest neighbors is generated by
@@
The law $\mathbb{P}$ of the environment is assumed to be stationary and ergodic with respect to the shift $\{\theta_{x}\}$ on $\mathbb{Z}^{2}$. For fixed environment $\omega$, we denote by $P^{\omega}_{0}$ the quenched law of the walk with $P^{\omega}_{0}(X(0)=0)=1.$ We are concerned with the question whether, for $\mathbb{P}$ a.e. $\omega$, under $P^{\omega}_{0}$, no explosion occurs and, whether the quenched invariance principle (QIP) holds, that is the diffusively rescaled walk
$X^{(T)}(t)\equiv T^{-1/2}X(Tt),\qquad t\geq 0$
converges in law as $T\to\infty$ to a non degenerate Brownian motion
@@
is sufficient for a QIP to hold. It should be noted, that in case of i.i.d. conductances defined on $\mathbb{Z}^{d}$, with $d\geq 2$, no such moment condition is required for the QIP, cf. [2], however for general ergodic $\mathbb{P}$ the moment condition is relevant for both non explosion and QIP. Generalizations of the results of [4] to environments that are non-symmetric have been obtained in the recent article [16].
@@
where $e_{1}=(1,0)$ and $e_{2}=(0,1)$, and where $(H(k))_{k\in\mathbb{Z}}$ and $(V(k))_{k\in\mathbb{Z}}$ are two independent families of i.i.d. positive random variables. Of course this implies the ergodicity of $\mathbb{P}$, which is however not directional ergodic.
@@
$X^{\tau_{R}}_{1}>(t)=2A^{H}_{1}(X_{2})(t\wedge\tau_{R})=\int_{0}^{t\wedge\tau_{R}}2H(X_{2}(s))\,ds,$
$X^{\tau_{R}}_{2}>(t)=2A^{V}_{2}(X_{1})(t\wedge\tau_{R})=\int_{0}^{t\wedge\tau_{R}}2V(X_{1}(s))\,ds,$
$(<X^{\tau_{R}}_{1},X^{\tau_{R}}_{2}>(t)=0)$. The additive functionals above can be interpreted as random walks in random scenery, although here the random walk is dependent of the ambient scenery.
Regarding the jump rates, we assume that the laws of $\{H(x_{2}),x_{2}\in\mathbb{Z}\}$ and $\{V(x_{1}),x_{1}\in\mathbb{Z}\}$ are i.i.d. with tails given, as $L\to\infty$, by
@@
The case 0 where $V(x_{1})=\text{const}$ was the object of 2 joint papers [14] and [15], where heat kernel estimates were derived.
In neither cases 1,2 it is a priori clear whether non explosion occurs. Also the random walk might travel very far along lines with very large conductances, thus we expect a super diffusive behaviour, i.e. a singular scaling.
@@
Moreover assume that for some $\gamma_{i}\geq 1/2$
@@
$\epsilon_{1}=\frac{{1}/{\alpha_{1}}-1}{2}\vee 0,\qquad\epsilon_{2}=\frac{{1}/{\alpha_{2}}-1}{2}\vee 0$
@@
(needed in order to get $\gamma_{i}>0$), we have
$\gamma_{1}=\frac{1+\epsilon_{1}}{2(1-\epsilon_{1}\cdot\epsilon_{2})}\geq 1/2\qquad\gamma_{2}=\frac{1+\epsilon_{2}}{2(1-\epsilon_{1}\cdot\epsilon_{2})}\geq 1/2.$
@@
we expect that when $\epsilon_{1}\cdot\epsilon_{2}>1$ explosion occurs.
Moreover with $\gamma_{i}^{+}>\gamma_{i}\geq 1/2$ and
@@
we show that for $\mathbb{P}$ a.e. $\omega$, for all $T>0$, we have the convergence
$\lim_{\epsilon\to 0}\sup_{0\leq t\leq T}|X_{i}^{T,+}(t)|=0,$
in $P^{\omega}_{0}$-probability, see Proposition 3.4 below.
In case 2 we can be more precise: in that case since $\epsilon_{2}=0$, no explosion occurs for all $\alpha_{1}>0$, and with
@@
then $\mathbb{P}$ a.s., as $\epsilon\to 0$, the law of
$X^{(T)}(t)=(T^{-\gamma_{1}}X_{1}(Tt),T^{-1/2}X_{2}(Tt)),\qquad t\geq 0$
under $P^{\omega}_{0}$ converges to
@@
where $B_{i}$ are independent Brownian motions and
$\Delta^{\mathcal{H}}(B_{2})(t)=\int_{\mathbb{R}}L^{B_{2}}(t,x)d\mathcal{H}(x),$
where $(\mathcal{H}(x))_{x\in\mathbb{R}}$ is a two-sided stable subordinator of index $\alpha_{1}$ and, for all $(t,x)\in\mathbb{R}_{+}\times\mathbb{R}$, $L^{B_{2}}(t,x)$ is the local time of $B_{2}$, up to time $t$, at the point $x$. See Theorem 4.3 below. Processes such as $\Delta^{\mathcal{H}}(B_{2})$ appear in the work of Kesten and Spitzer [19] to describe the scaling limits of the models of random walk in random scenery considered therein.
Since in case 2, $\mathbb{E}[V(x_{1})]<\infty$, the convergence of $X_{2}^{(T)}(t)\equiv T^{-1/2}X_{2}(Tt)$ to $B_{2}(t)$ follows from a martingale scaling limit theorem, however the proof of the convergence of the local times, based on an appropriate time change argument, is not immediate. In case 1 we expect that $X^{(T)}$ converges in law, for the Skorohod topology, to a process $(Z_{1},Z_{2})$ of the form
@@
Besides the line model random walk considered above, we also consider in this work the associated constant speed random walk $(\hat{X}(t))_{t\geq 0}$ generated by
@@
which corresponds to a simple time change of the original walk $X$, namely we can construct $\hat{X}$ as $\hat{X}(t)=X\circ J^{H,V}(X)^{-1}(t)$, where
$J^{H,V}(X)^{-1}(t)=\inf\{s>0:J^{H,V}(X)(s)\geq t\},\qquad t\geq 0,$
is the inverse of the increasing function $J^{H,V}(X)(s):=[A^{H}_{1}(X_{2})+A^{V}_{2}(X_{1})](s)$, $s\geq 0$.
In case 2, the random walk is trapped along the vertical direction and $\hat{X}_{2}$ sub-diffusively rescaled converges to a FIN diffusion as considered in [17] and [6], whereas in the horizontal direction $\hat{X}_{1}$ rescales diffusively. More precisely with $\gamma_{1}=\frac{1}{2}(\frac{1}{2\alpha_{1}}+\frac{1}{2})$, (so that we have $1/(4\gamma_{1})<1/2$), set
$\hat{X}^{(T)}(t)=(T^{-1/2}\hat{X}_{1}(t/\epsilon),T^{-1/(4\gamma_{1})}\hat{X}_{2}(Tt))$
then for $\mathbb{P}$ a.e. $\omega$, under $P^{\omega}_{0}$, $(\hat{X}^{(T)}(t))_{t\geq 0}$ converges in law to $(B_{1}(t),B_{2}\circ\Delta^{\mathcal{H}}(B_{2})^{-1}(t)$ where $B_{i},i=1,2$, are independent Brownian motions and
@@
is the inverse of the process $\Delta^{\mathcal{H}}(B_{2})(s)=\int_{\mathbb{R}}L^{B_{2}}(s,x)d\mathcal{H}(x)$, $s\geq 0$, and where $(\mathcal{H}(x))_{x\in\mathbb{R}}$ is again a two-sided stable subordinator of index $\alpha_{1}$, see Theorem 4.11 below.
@@
$\hat{X}^{(T)}(t)=(T^{-1/2}\hat{X}_{1}(Tt),T^{-\gamma_{2}/(2\gamma_{1})}\hat{X}_{2}(Tt)),$
converges to $(B_{1}(t),B_{2}\circ(\Delta^{\mathcal{V}}(Z_{1})\circ\Delta^{\mathcal{H}}(Z_{2})^{-1}(t)))$, where $\Delta^{\mathcal{V}}(Z_{1})$ and $\Delta^{\mathcal{H}}(Z_{2})$ are as in (1.4) and
@@ 2.1 Ergodicity of the environment
Under the annealed law $\mathbf{P}=\int\mathbb{P}(d\omega)\,P^{\omega}_{0}$, this defines a Markov process on $\Omega$, with reversible measure $\mathbb{P}$. By ergodicity of the environment, this process is furthermore ergodic, see e.g. [25, Corollary 2.1.25].
@@ 2.2 Martingale property of the random walk
$X^{\tau_{R}}_{1}>(t)=2A^{H}_{1}(X_{2})(t\wedge\tau_{R}),\quad<X^{\tau_{R}}_{2}>(t)=2A^{V}_{2}(X_{1})(t\wedge\tau_{R}),\qquad t\geq 0.$
@@ 3 Non-explosion
Since $\varepsilon_{1}^{+},\varepsilon_{2}^{+}<1$, we get that $h_{2}(t),v_{2}(t)<\infty$ for all $t\geq 0$. More specifically, by the first point in Lemma 5.5 below (see subsection 5.3 in the Appendix) with $h_{0}=v_{0}=T$, we obtain
@@
Combining (3.8) and (3.9) we deduce, by the second point of Lemma 5.5 applied with $h_{0}=v_{0}=T$, that for all $t\in[0,T]$,
@@ 4.2.1 The result
$\left(\frac{1}{T^{\delta}}\int_{0}^{Tt}H^{(T)}(X_{2}(s))\,ds\right)_{t\geq 0}\underset{T\to\infty}{\Longrightarrow}\left(\Delta(t)\right)_{t\geq 0},$ (4.4)
@@ 4.2.2 Case where $V$ is constant
In this case, $X(t)=(X_{1}(t),X_{2}(t))$ can be represented as $(S_{1}(A(t)),S_{2}(t))$, where $S(t)=(S^{1}(t),S^{2}(t))$ is a 2-dimensional random walk jumping to each neighboring site at rate 1, and
@@
$\left(\frac{1}{T^{\delta}}\int_{0}^{Tt}H^{(T)}(X_{2}(s))\,ds\right)_{t\geq 0}\underset{T\to\infty}{\Longrightarrow}(\Delta(t))_{t\geq 0},$
@@
$\begin{split}&\frac{1}{T^{\delta}}\int_{0}^{Tt}H^{(T)}(X_{2}(r))\,dr=\int_{\mathbb{R}}\ell^{T}_{t}(x)\,\mu_{T}(dx)\\ =&\int_{\mathbb{R}}\varphi_{M}(x)\ell^{T}_{t}(x)\,\mu_{T}(dx)+\int_{\mathbb{R}}(1-\varphi_{M}(x))\ell^{T}_{t}(x)\,\mu_{T}(dx),\end{split}$ (4.8)

### v2

@@ Abstract
We consider a 2-dimensional model of random walk in random environment known as line model. The environment is described by two independent families of i.i.d. random variables dictating rates of jumps in vertical, respectively horizontal directions, and whose values are constant along vertical, respect. horizontal lines. When jump rates are heavy-tailed in one of the directions, and given by a stable distribution, we prove that the random walk becomes superdiffusive in that direction, with an explicit scaling limit written as a two-dimensional Brownian motion time-changed (in one of the components) by a Kesten-Spitzer process as introduced in [19]. In the case of a fully degenerate environment, a sufficient condition for non-explosion is provided, and conjectures on the associated scaling limit are formulated.
@@ 1.1 Outline of the article
This paper deals with a time-continuous random walk $(X(t))_{t\geq 0}$ on $\mathbb{Z}^{2}$ in the random environment $\omega$. The walk to nearest neighbors is generated by the (random) operator $\mathcal{L}^{\omega}$ acting on test functions $f:\mathbb{Z}^{2}\to\mathbb{R}$ as
@@
The law $\mathbb{P}$ of the environment is assumed to be stationary and ergodic with respect to the shifts $\{\theta_{x}\}$ on $\mathbb{Z}^{2}$. For fixed environment $\omega$, we denote by $P^{\omega}_{0}$ the quenched law of the walk with $P^{\omega}_{0}(X(0)=0)=1.$ We are concerned with the question whether, for $\mathbb{P}$ a.e. $\omega$, under $P^{\omega}_{0}$, no explosion occurs and, whether the quenched invariance principle (QIP) holds, that is the diffusively rescaled walk
$X^{(T)}(t)\equiv T^{-1/2}X(Tt),\qquad t\geq 0,$
converges in law as $T\to\infty$ to a non degenerate Brownian motion.
@@
is sufficient for a QIP to hold. It should be noted, that in case of i.i.d. conductances defined on $\mathbb{Z}^{d}$, with $d\geq 2$, no such moment condition is required for the QIP, cf. [2], however for general ergodic $\mathbb{P}$ the moment condition is relevant for both non explosion and QIP. Generalizations of the results of [4] to environments that are non-symmetric have been obtained in the recent article [16]. We also refer the reader to [20, Chap. 3], which establishes central limit theorems in $L^{1}$ with respect to the environment.
@@
for $x=(x_{1},x_{2})\in\mathbb{Z}^{2}$, where $e_{1}=(1,0)$ and $e_{2}=(0,1)$, and where $(H(k))_{k\in\mathbb{Z}}$ and $(V(k))_{k\in\mathbb{Z}}$ are two independent families of i.i.d. positive random variables. Of course this implies the ergodicity of $\mathbb{P}$, which is however not directional ergodic.
@@
$X^{\tau_{R}}_{1}>(t)=\int_{0}^{t\wedge\tau_{R}}2H(X_{2}(s))\,ds,\qquad<X^{\tau_{R}}_{2}>(t)=\int_{0}^{t\wedge\tau_{R}}2V(X_{1}(s))\,ds,$
while we have $X^{\tau_{R}}_{1},X^{\tau_{R}}_{2}>(t)=0$ for all $t\geq 0$. The additive functionals above can be interpreted as random walks in random scenery, although here the random walk is dependent of the ambient scenery.
Regarding the jump rates, we assume $\{H(x_{2}),x_{2}\in\mathbb{Z}\}$ and $\{V(x_{1}),x_{1}\in\mathbb{Z}\}$ to be two independent families of i.i.d. positive random variables, with tails given, as $L\to\infty$, by
@@
The case where the random variables $V(x_{1})$ are a deterministic constant was the object of 2 joint papers [14] and [15], where heat kernel estimates were derived.
In neither cases 1,2 is it a priori clear whether non explosion occurs. Also the random walk might travel very far along lines with very large conductances, thus we expect a super diffusive behaviour, i.e. a singular scaling.
@@
Moreover assume that, for some $\gamma_{i}\geq 1/2$,
@@
$\epsilon_{1}=\frac{{1}/{\alpha_{1}}-1}{2}\vee 0,\qquad\epsilon_{2}=\frac{{1}/{\alpha_{2}}-1}{2}\vee 0,$
@@
(needed in order to get $\gamma_{i}>0$), we get
$\gamma_{1}=\frac{1+\epsilon_{1}}{2(1-\epsilon_{1}\cdot\epsilon_{2})}\geq 1/2,\qquad\gamma_{2}=\frac{1+\epsilon_{2}}{2(1-\epsilon_{1}\cdot\epsilon_{2})}\geq 1/2.$
@@
we expect that when $\epsilon_{1}\cdot\epsilon_{2}>1$ explosion occurs. Moreover with $\gamma_{i}^{+}>\gamma_{i}\geq 1/2$ and
@@
we show that for $\mathbb{P}$ a.e. $\omega$, for all $t>0$, the convergence
$\lim_{T\to\infty}\sup_{0\leq s\leq t}|X_{i}^{T,+}(s)|=0,\qquad i=1,2,$
holds in $P^{\omega}_{0}$-probability, see Proposition 3.4 below.
In case 2 we can be more precise: in that case since $\epsilon_{2}=0$, no explosion occurs for all $\alpha_{1}>0$. Moreover, assuming that the random variables $H(x)$, $x\in\mathbb{Z}$, follow a stable distribution of index $\alpha_{1}$, then setting
@@
and performing an appropriate rescaling of the environment together with the random walk, then for $\mathbb{P}$-a.e. realisation of the environment, as $T\to\infty$, the process
$X^{(T)}(t)=(T^{-\gamma_{1}}X_{1}(Tt),T^{-1/2}X_{2}(Tt)),\qquad t\geq 0,$
converges in law, and for the Skorohod topology, to the process
@@
where $B_{1},B_{2}$ are two independent Brownian motions and
$\Delta^{\mathcal{H}}(B_{2})(t)=\int_{\mathbb{R}}L^{B_{2}}(t,x)d\mathcal{H}(x),\qquad t\geq 0.$ (1.3)
Here $(\mathcal{H}(x))_{x\in\mathbb{R}}$ is a two-sided stable subordinator of index $\alpha_{1}$ and, for all $(t,x)\in\mathbb{R}_{+}\times\mathbb{R}$, $L^{B_{2}}(t,x)$ is the local time of $B_{2}$, up to time $t$, at the point $x$. See Theorem 4.3 below for the precise statement.
Processes such as $\Delta^{\mathcal{H}}(B_{2})$ appear in the work of Kesten and Spitzer [19] to describe the scaling limits of the models of random walk in random scenery considered therein. Since in case 2, $\mathbb{E}[V(x_{1})]<\infty$, the convergence of $X_{2}^{(T)}(t)\equiv T^{-1/2}X_{2}(Tt)$ to $B_{2}(t)$ follows from a martingale scaling limit theorem, wherefrom the results of [19] heuristically entail the convergence of the first additive functional $A^{H}_{1}(t)$ of (1.1), when properly rescaled, to (1.3). However, in contrast to [19] a major caveat arises in our setting, which is the lack of independence of the process $(X_{2}(t))_{t\geq 0}$ with the random variables $H(k)$, $k\in\mathbb{Z}$. In addition, the convergence of the local times of $X_{2}^{(T)}$ towards those of $B_{2}$ is not clear at all. To overcome these difficulties, we resort to an appropriate time-change, which transforms the diffusive component $X_{2}$ into a simple random walk. Undoing the time change hinges on an appropriate application of Birkhoff’s ergodic theorem for the environment seen from the walker and relies on the random variables $V(k)$, $k\in\mathbb{Z}$, being integrable.
In case 1 we expect that $X^{(T)}$ converges in law, for the Skorohod topology, to a process $(Z_{1},Z_{2})$ of the form
@@
Besides the (variable-speed) line model random walk $(X(t))_{t\geq 0}$ considered above, we also consider in this work the associated constant speed random walk $(\hat{X}(t))_{t\geq 0}$ generated by
@@
which corresponds to a simple time change of the original walk $X$, namely we can construct $\hat{X}$ as $\hat{X}(t)=X\circ{J^{H,V}}^{-1}(t)$, where
${J^{H,V}}^{-1}(t)=\inf\{s>0:J^{H,V}(s)\geq t\},\qquad t\geq 0,$
is the inverse of the increasing function $J^{H,V}(s):=A^{H}_{1}(s)+A^{V}_{2}(s)$, $s\geq 0$ (see (1.1)).
In case 2, assuming again the random variables $H(k)$, $k\in\mathbb{Z}$, to follow an $\alpha_{1}$-stable distribution, then the random walk is trapped along the vertical direction and $\hat{X}_{2}$ sub-diffusively rescaled converges to a FIN diffusion as considered in [17] and [6], whereas in the horizontal direction $\hat{X}_{1}$ rescales diffusively. More precisely with $\gamma_{1}=\frac{1}{2}(\frac{1}{2\alpha_{1}}+\frac{1}{2})$, (so that we have $1/(4\gamma_{1})<1/2$), set
$\hat{X}^{(T)}(t)=(T^{-1/2}\hat{X}_{1}(t/\epsilon),T^{-1/(4\gamma_{1})}\hat{X}_{2}(Tt)),\qquad t\geq 0.$
Then, performing an appropriate rescaling of the environment together with the random walk, for $\mathbb{P}$ a.e. $\omega$, the process $(\hat{X}^{(T)}(t))_{t\geq 0}$ converges in law for the Skorohod topology, as $T\to\infty$, to $(B_{1}(t),B_{2}\circ\Delta^{\mathcal{H}}(B_{2})^{-1}(t))$ where $B_{1},B_{2}$ are two independent Brownian motions and
@@
is the inverse of the process $\Delta^{\mathcal{H}}(B_{2})(s)=\int_{\mathbb{R}}L^{B_{2}}(s,x)d\mathcal{H}(x)$, $s\geq 0$, with $(\mathcal{H}(x))_{x\in\mathbb{R}}$ a two-sided stable subordinator of index $\alpha_{1}$, see Theorem 4.11 below.
@@
$\hat{X}^{(T)}(t)=(T^{-1/2}\hat{X}_{1}(Tt),T^{-\gamma_{2}/(2\gamma_{1})}\hat{X}_{2}(Tt)),\qquad t\geq 0,$
converges as $T\to\infty$ to $(B_{1}(t),B_{2}\circ(\Delta^{\mathcal{V}}(Z_{1})\circ\Delta^{\mathcal{H}}(Z_{2})^{-1}(t)))$, where $\Delta^{\mathcal{V}}(Z_{1})$ and $\Delta^{\mathcal{H}}(Z_{2})$ are as in (1.5) and
@@ 2.1 Ergodicity of the environment
Under the annealed law $\mathbf{P}=\int\mathbb{P}(d\omega)\,P^{\omega}_{0}$, this defines a Markov process on $\Omega$, with reversible measure $\mathbb{P}$. By ergodicity of the environment, this process is furthermore ergodic, see e.g. [26, Corollary 2.1.25].
@@ 2.2 Martingale property of the random walk
$X^{\tau_{R}}_{1}>(t)=2A^{H}_{1}(t\wedge\tau_{R}),\quad<X^{\tau_{R}}_{2}>(t)=2A^{V}_{2}(t\wedge\tau_{R}),\qquad t\geq 0.$
@@ 3 Non-explosion
Since $\varepsilon_{1}^{+},\varepsilon_{2}^{+}<1$, we get that $h_{2}(t),v_{2}(t)<\infty$ for all $t\geq 0$. More specifically, by the first point in Lemma 5.5 below (see subsection 5.3 in the Appendix) with $\kappa=T$, we obtain
@@
Combining (3.8) and (3.9) we deduce, by the second point of Lemma 5.5 applied with $\kappa=T$, that for all $t\in[0,T]$,
@@ 4.2.1 The result
$\left(\frac{1}{T^{\delta}}\int_{0}^{Tt}H^{(T)}(X^{T}_{2}(s))\,ds\right)_{t\geq 0}\underset{T\to\infty}{\Longrightarrow}\left(\Delta(t)\right)_{t\geq 0},$ (4.4)
@@ 4.2.2 Case where $V$ is constant
In this case, $X(t)=(X_{1}(t),X_{2}(t))$ can be represented as $(S_{1}(A(t)),S_{2}(t))$, where $S(t)=(S_{1}(t),S_{2}(t))$ is a 2-dimensional random walk jumping to each neighboring site at rate 1, and
@@
$\left(\frac{1}{T^{\delta}}\int_{0}^{Tt}H^{(T)}(S_{2}(r))\,dr\right)_{t\geq 0}\underset{T\to\infty}{\Longrightarrow}(\Delta(t))_{t\geq 0},$
@@
$\begin{split}&\frac{1}{T^{\delta}}\int_{0}^{Tt}H^{(T)}(S_{2}(r))\,dr=\int_{\mathbb{R}}\ell^{T}_{t}(x)\,\mu_{T}(dx)\\ =&\int_{\mathbb{R}}\varphi_{M}(x)\ell^{T}_{t}(x)\,\mu_{T}(dx)+\int_{\mathbb{R}}(1-\varphi_{M}(x))\ell^{T}_{t}(x)\,\mu_{T}(dx),\end{split}$ (4.8)